Grade upto college level Algebra

The differential of the family of circles with fixed radius 5 units and centre on the line y = 2 is

aditya kashyap
12 Years agoGrade upto college level

1 Answer

Jitender Singh
12 Years ago
Ans:
Equation of Circle:
(x-a)^{2}+(y-2)^{2} = (5)^{2}…............(1)
Differentiate:
2(x-a)+2(y-2).y^{'} = 0
(x-a) =-(y-2).y^{'}
Put this value in (1)
(y-2)^{2}(1+(y^{'})^{2}) = 25
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty